OpenStudy (ginabean819):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 123 days and there is an amount equal to 150 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 365 days? Hint: The exponential equation for half-life is A(t) = A0(0.5)t/H, where A(t) is the final amount remaining, A0 is the initial amount, t is time, and H is the half-life.

OpenStudy (ginabean819):

A(t) = 150(0.5)t/123, 19.2 grams remaining A(t) = 150(0.5)123t, 1.4 grams remaining A(t) = 123(150)(0.5)t, 0.0 gram remaining A(t) = 150(0.5)123/t, 118.8 grams remaining

OpenStudy (wolf1728):

That equation needs to be made a little bit easier to read: (and Formula Attached). ending amount = (beginning amt)*[.5^(time/half-life)] ending amount = (150)*[.5^(365/123)] ending amount = (150)*[.5^2.9674796748] I think you can finish it from there.